OpenStudy (love_to_love_you):

Write an equation in point slope form, that passes through points (3, -10) and (-5, -2). a) y-10=x-3 b) y+10=-(x-3) c) y+10=2(x-3) d) y-10=1/2(x-3)

jigglypuff314 (jigglypuff314):

point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1)

OpenStudy (anonymous):

You need to find the slope first. Use \[m=\frac{ x2-x1 }{ y2-y1 }\]

OpenStudy (love_to_love_you):

So we can eliminate a and d

OpenStudy (love_to_love_you):

slope is -1

OpenStudy (anonymous):

The slope is -1

OpenStudy (love_to_love_you):

so b?

OpenStudy (anonymous):

Yes.

jigglypuff314 (jigglypuff314):

yes I think B is correct ^_^

OpenStudy (anonymous):

i mean b sorry i am sooo busy

OpenStudy (love_to_love_you):

Okay, thanks y'all!

OpenStudy (anonymous):

:D